# Higher-Order Message Passing for Glycan Representation Learning

### v1

@@ Appendix B Further Ablation Studies
Figure A2B compares different pooling mechanisms for the GIFFLAR model with 128-dimensional feature vectors. Here, we only compared the models on the Glycosylation, Immunogenicity, and Domain datasets to reduce runtime. The six modes we compared were:
@@
Surprisingly, attention-free pooling functions generally performed better than attention-based ones in this analysis. Further, global poolings were better than local or weighted local methods. We conclude that global mean pooling seems to constitute the most performant pooling operation for GIFFLAR.
Appendix C Model Comparison
@@
We report Matthews Correlation Coefficients (MCC) for the final models and baselines to give more perspective on the model performances (Table A1). For the multi-label datasets, the MCC was computed per label and then averaged over the labels unweighted.
@@
Appendix D Training Details
Table A2 lists some core measurements of the models we compare. We used the Domain prediction from the taxonomy collection as an example dataset. The table’s average number of nodes per graph is computed on the training set. The first three models (RF, SVM, and XGB) were trained using scikit-learn v1.5.1 [31] on a regular CPU. The deep learning models were trained using PyTorch v2.3.1 [32], PyTorch Geometric v2.5.3 [33], and PyTorch Lightning v2.3.2 [34] on an NVIDIA GeForce RTX 3090 with 24GB GPU RAM. All metrics for all models were computed using TorchMetrics v1.4.0 [35].
The code and the data splits are available on GitHub link-removed-for-double-blind-review.
Model Training time $\varnothing$ # nodes per graph # param.
Morgan Fingerprint (1024-bit)
RF 0.1 min
SVM 2.2 min
XGB 1.7 min
MLP 11.7 min 527K
Homogeneous Graphs
GNNGLY 20.5 min 91 8M
SweetNet 22.7 min 13 37.2M
GLAMOUR 9.5 min 7 18.2M
Heterogeneous Graphs
$\text{RGCN}_{\text{RW}}$ 55.3 min 195 41.9M
GIFFLAR 45.0 min 195 35.1M

### v3

@@ 5.1 Ablation Studies
Figure 2d compares different pooling mechanisms for the GIFFLAR model with 128-dimensional feature vectors. Here, we only compared the models on the Glycosylation, Immunogenicity, and Domain datasets to reduce runtime. The six modes we compared were:
@@
Surprisingly, attention-free pooling functions generally performed better than attention-based ones in this analysis. Further, global pooling was better than local or weighted local methods. We conclude that global mean pooling seems to constitute the most performant pooling operation for GIFFLAR.
We also conducted ablation studies to determine the backbone structure of GLAMOUR and the influence of PEs on the performance of RGCN. The results are described in Appendix B.
6 Discussion, Conclusion, Limitations, and Future Work
Here, we presented GIFFLAR, a novel graph neural network architecture designed specifically for glycan representation learning. By leveraging combinatorial complexes and higher-order message passing, GIFFLAR achieves state-of-the-art performance across a wide range of glycan property prediction tasks. GIFFLAR consistently outperforms existing methods, including traditional machine learning approaches using Morgan fingerprints, homogeneous graph neural networks such as GNNGLY, SweetNet, and GLAMOUR, and heterogeneous graph models such as RGCN. The superior performance of GIFFLAR can likely be attributed to its ability to capture both atomic-level structure and topological information of glycans simultaneously. Due to its superior performance, we expect that GIFFLAR will become the new foundation in computational glycobiology and yield improvements in relevant downstream tasks. Another advantage of GIFFLAR over the baselines is that the feature extractor can be trained end-to-end specifically for a task, e. g., in lectin-glycan interaction prediction.
In conclusion, GIFFLAR presents a significant advance in glycan representation learning, demonstrating the power of combining combinatorial complexes with higher-order message passing in graph neural networks. As glycomics continues to grow in importance within the life sciences, we envision that GIFFLAR and its future extensions will play a crucial role in unlocking the full potential of glycan-related research and applications. A prime example application of GIFFLAR as a trainable feature extractor for glycans would be for instance the field of lectin-glycan interaction prediction.
Acknowledgments and Disclosure of Funding
R.J. was supported by the Knut and Alice Wallenberg Foundation and the University of Gothenburg. This work was funded by a Branco Weiss Fellowship - Society in Science awarded to D.B., the Knut and Alice Wallenberg Foundation, and the University of Gothenburg, Sweden.
The authors thank the Drug Bioinformatics group at the Helmholtz Institute for Pharmaceutical Research Saarland, Saarbrücken, Germany, for providing the computing resources to conduct the experiments.
Code Availability
The implementation of GIFFLAR is publicly available at github.com/BojarLab/GIFFLAR.
Data Availability
The data (including the data splitting) used in this study is publicly available in the datasets folder of github.com/BojarLab/GIFFLAR.
Author Contributions
R.J. implemented the code and conducted the experiments. R.J. and D.B. wrote the manuscript.
Competing Interests
R.J. has no competing interests. D.B. consults for SweetSense Analytics AB.
Appendix A Training Details
Table A1 lists core measurements of the models we compare. We used the Domain prediction from the taxonomy collection as an example dataset. The table’s average number of nodes per graph is computed on the training set. The first three models (RF, SVM, and XGB) were trained using scikit-learn v1.5.1 [40] on a regular CPU. The deep learning models were trained using PyTorch v2.3.1 [41], PyTorch Geometric v2.5.3 [42], and PyTorch Lightning v2.3.2 [43] on an NVIDIA GeForce RTX 3090 with 24GB GPU RAM. All metrics for all models were computed using TorchMetrics v1.4.0 [44].
Model Training time $\varnothing$ # nodes per graph # param.
Morgan Fingerprint (1024-bit)
RF 0.1 min
SVM 2.2 min
XGB 1.7 min
MLP 11.7 min 527K
Homogeneous Graphs
GNNGLY 20.5 min 91 8M
SweetNet 22.7 min 13 37.2M
GLAMOUR 9.5 min 7 18.2M
Heterogeneous Graphs
$\text{RGCN}_{\text{RW}}$ 55.3 min 195 41.9M
GIFFLAR 45.0 min 195 35.1M
Appendix B Further Model Analysis
To investigate the impact of Algorithm 1 on the performance comparison in Figure 2, we report the same plot but unnormalized in Figure A1 (except for MCC, which was transformed by $x\mapsto\frac{x+1}{2}$ to be in the interval [0,1]). Compared to Figure 2, the ordering of the models by performance does not change significantly, but the differences between the models shrink.
Figure A1: a-d Reevaluation of Figure 2 without the normalization in Algorithm 1. b contains additional results for testing PEs for the RGCN model. e Comparison of three backbones for the GLAMOUR model.
@@ Appendix B Further Model Analysis
We report Matthews Correlation Coefficients (MCC) for the final models and baselines to give more perspective on the model performances (see Table A2). For the multi-label datasets, the MCC was computed per label and then averaged over the labels unweighted.
As Mohapatra et al. do not name one backbone architecture as the best among the five provided [19], we compared the usable ones. The provided implementation does not allow for applying GCN— and GAT-based models to graphs representing monomeric molecules. In Figure A1a, we compared the Weave, MPNN, and AttentiveFP models, among which the MPNN performed best and was selected as the GLAMOUR backbone for the baseline model in this work.
As the RGCN model is the only baseline without specified featurization of the nodes, we tested PEs here as well and found that adding positional encodings from random walks improved the performance as shown in Figure A1b.